Explain how 14/20 is not in simplest form.
Answer by stanbon(75887)

Explain how 14/20 is not in simplest form.
----------
14/20 = (7*2)/(10*2) = (7/10)(2/2) = (7/10)(1) = 7/10
---------------------------------------------
Cheers,
Stan H.
-----------------